# Predictively Oriented Posteriors

### v1

@@ Appendix C Proofs of Main results
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q,k)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}_{r}(\theta),P_{0}\right)+{\mathsf{AppGap}}(Q^{\star}_{r},k)+\frac{\textnormal{KL}(Q^{\star}_{r}\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$=\inf_{r>0}\left\{\mathcal{D}_{S}\left(\sum_{j=1}^{K}\omega_{j}\int P_{\theta}\,\mathsf{d}\pi^{\star}_{r,j}(\theta),\sum_{j=1}^{K}\omega_{j}P_{\theta_{j}^{\star}}\right)+\frac{\textnormal{KL}\left(\sum_{j=1}^{K}\omega_{j}\Pi^{\star}_{r,j}\|\Pi\right)}{\lambda_{n}}+{\mathsf{AppGap}}(Q^{\star}_{r},k)\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\pi^{\star}_{r,j}(\theta)+\frac{\sum_{j=1}^{K}\omega_{j}\textnormal{KL}\left(\Pi^{\star}_{r,j}\|\Pi\right)}{\lambda_{n}}+{\mathsf{AppGap}}(Q^{\star}_{r},k)\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\pi^{\star}_{r,j}(\theta)+\frac{\sum_{j=1}^{K}\omega_{j}\log\{1/\Pi(\mathcal{B}_{r,j})\}}{\lambda_{n}}+{\mathsf{AppGap}}(Q^{\star}_{r},k)\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\pi^{\star}_{r,j}(\theta)+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}+{\mathsf{AppGap}}(Q^{\star}_{r},k)\right\}+\frac{\lambda_{n}C_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{r+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}\right\}+{\mathsf{AppGap}}(Q^{\star}_{r},k)+\frac{\lambda_{n}C_{k}^{2}}{n}.$
Optimising $r$ using similar arguments to the end of Lemma 5 then yields the stated result. Furthermore, when equation (6) is satisfied with equality, ${\mathsf{AppGap}}(Q^{\star}_{r},k)=0$ for any $r>0$. ∎
@@
The stated claim follows from the results in Theorem 3 and Lemma 5. In particular, from the proof of Theorem 3, we have,
$\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]\leq\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+\inf_{r>0}\left\{r+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}\right\}+\frac{\lambda_{n}C_{k}^{2}}{n},$
and optimizing the choice of $\lambda_{n}$ as in the proof of Lemma 5 then yields
$0\leq\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\leq$ $\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})+C\frac{\log(n)}{n^{1/2}}$
$\leq$ $C\frac{\log(n)}{n^{1/2}}+\mathbb{E}[\mathcal{D}_{S}(P_{Q^{\dagger}_{n}},P_{0})]+\mathbb{E}\left\{\mathcal{D}_{S}(P_{\theta^{\star}},P_{0})-\mathcal{D}_{S}(P_{Q^{\dagger}_{n}},P_{0})\right\}$
$\leq$ $C\frac{\log(n)}{n^{1/2}}+\mathbb{E}[\mathcal{D}_{S}(P_{Q^{\dagger}_{n}},P_{0})],$
where the third line follows from Lemma 5. ∎
@@
Since $\mathsf{d}_{2}(\cdot,\cdot)$ satisfies a weak triangle inequality, for some $C>0$:
                                                                                                    $\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{n}^{\dagger}})]$ ${}\leq{}$ $C\mathbb{E}\{\mathsf{d}_{2}(P_{Q_{n}},P_{0})+\mathsf{d}_{2}(P_{Q_{n}^{\dagger}},P_{0})\}.$                                                                                                      (35)
Recall that, by Theorem 5, $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]\leq C\log(n)/n^{1/2}$, while from Lemma 5, we also have $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\leq C\log(n)/n^{1/2}$. Hence, from Assumption 8(iii), and concavity of the map $x\mapsto x^{1/\alpha}$ for $\alpha\geq 1$,
$\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{0})]\leq$ $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})^{1/\alpha}]\leq\left\{\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}},P_{0})]\right\}^{1/\alpha}\leq C\{\log(n)/n^{1/2}\}^{1/\alpha},$
$\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}^{\dagger}},P_{0})]\leq$ $\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})^{1/\alpha}]\leq\left\{\mathbb{E}[\mathcal{D}_{S}(P_{Q_{n}^{\dagger}},P_{0})]\right\}^{1/\alpha}\leq\{\log(n)/n^{1/2}\}^{1/\alpha}.$
where the first bound comes from Theorem 5, and the second from Lemma 5. Applying the above bounds into (C), we obtain
                                                                                                    $\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{n}^{\dagger}})]$ ${}\leq{}$ $C\{\log(n)/n^{1/2}\}^{1/\alpha}.$                                                                                                      (36)

### v2

@@ Appendix C Proofs of Main results
$\leq\inf_{Q\in\mathcal{P}(\Theta)}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\mathsf{d}Q(\theta),P_{0}\right)+\frac{\text{KL}(Q\|\Pi)}{\lambda_{n}}\right\}+\frac{\lambda_{n}kC_{k}^{2}}{n}$
$\leq\inf_{r>0}\left\{\mathcal{D}_{S}\left(\int P_{\theta}\,\mathsf{d}Q^{\star}_{r}(\theta),P_{0}\right)+\frac{\textnormal{KL}(Q^{\star}_{r}\|\Pi)}{\lambda_{n}}\right\}+\nu_{n}$
$=\inf_{r>0}\left\{\mathcal{D}_{S}\left(\sum_{j=1}^{K}\omega_{j}\int P_{\theta}\,\mathsf{d}\Pi^{\star}_{r,j}(\theta),\sum_{j=1}^{K}\omega_{j}P_{\theta_{j}^{\star}}\right)+\frac{\textnormal{KL}\left(\sum_{j=1}^{K}\omega_{j}\Pi^{\star}_{r,j}\|\Pi\right)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\Pi^{\star}_{r,j}(\theta)+\frac{\sum_{j=1}^{K}\omega_{j}\textnormal{KL}\left(\Pi^{\star}_{r,j}\|\Pi\right)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\Pi^{\star}_{r,j}(\theta)+\frac{\sum_{j=1}^{K}\omega_{j}\log\{1/\Pi(\mathcal{B}_{r,j})\}}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{r>0}\left\{\sum_{j=1}^{K}\omega_{j}\,\int\,\mathcal{D}_{S}\left(P_{\theta},P_{\theta_{j}^{\star}}\right)\mathsf{d}\Pi^{\star}_{r,j}(\theta)+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}\right\}+\nu_{n}$
$\leq\inf_{r>0}\left\{r+\frac{d_{s}\log(c_{s}/r)}{\lambda_{n}}\right\}+\nu_{n}.$
Optimising $r$ using similar arguments to the end of Lemma 3 yields the stated result. ∎
@@
The proof follows directly from Theorem 4; in particular,
$\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]-\mathbb{E}[\mathcal{D}_{S}(P_{Q^{\dagger}_{n}},P_{0})]\leq\mathbb{E}\left[\mathcal{D}_{S}\left(P_{Q_{n}},P_{0}\right)\right]\lesssim\nu_{n}\,.$
@@
Since $\mathsf{d}_{2}(\cdot,\cdot)$ satisfies a weak triangle inequality, for some $C>0$,
                                                                                                    $\mathbb{E}[\mathsf{d}_{2}(P_{Q_{n}},P_{Q_{\star}})]$ ${}\leq{}$ $C\mathbb{E}\{\mathsf{d}_{2}(P_{Q_{n}},P_{0})+\mathsf{d}_{2}(P_{Q_{\star}},P_{0})\}$                                                                                                      (30)